Given the following national income model:

and


𝑇=𝑇 +π‘‘π‘Œ 0

π‘Œ = 𝐢 + 𝐼0 + 𝐺0

𝐢 = 𝐢0 + 𝑏(π‘Œ βˆ’ 𝑇)Β 


where π‘Œ = π‘–π‘›π‘π‘œπ‘šπ‘’, 𝐢 = π‘π‘œπ‘›π‘ π‘’π‘šπ‘π‘‘π‘–π‘œπ‘›, 𝑇 = π‘‘π‘Žπ‘₯π‘Žπ‘‘π‘–π‘œπ‘›;

𝑀hπ‘’π‘Ÿπ‘’ 𝐢0 > 0; 0 < 𝑏 < 1; 0 < 𝑑 < 1

Using Cramer’s rule to derive an expression for the equilibrium level of income, consumption, and taxation.Β 


1
Expert's answer
2021-05-04T12:15:06-0400

Solution:

First re-arrange the equations so that the endogenous variables are on the LHS and the exogenous variables are on the RHS:

Y βˆ’ C = I + G0

C βˆ’ bY + bT = C0

T – tY0 = T

Β 

Then rewrite the equations in matrix form:


The equilibrium national income (Y) = C + I0 + G0

Y = C0 + b(Y – T) + I0Β +Β G0


"=\frac{C0 − bT + I0 + G0
}{1 − b(1 − t)}"


Consumption:


"=\frac{C0 − bT + b (1 − t) (I0 + G0)

}{1 − b(1 − t)}"


Taxation:


"==\frac{C0 − bT + at + t(I0 + G0))

}{1 − b(1 − t)
}"